# Neural Guided Diffusion Bridges

### v2

@@ Abstract
We propose a novel method for simulating conditioned diffusion processes (diffusion bridges) in Euclidean spaces. By training a neural network to approximate bridge dynamics, our approach eliminates the need for computationally intensive Markov Chain Monte Carlo (MCMC) methods or reverse-process modeling. Compared to existing methods, it offers greater robustness across various diffusion specifications and conditioning scenarios. This applies in particular to rare events and multimodal distributions, which pose challenges for score-learning- and MCMC-based approaches. We propose a flexible variational family for approximating the diffusion bridge path measure which is partially specified by a neural network. Once trained, it enables efficient independent sampling at a cost comparable to sampling the unconditioned (forward) process.
@@ 1 Introduction
Diffusion processes play a fundamental role in various fields such as mathematics, physics, evolutionary biology, and, recently, generative models. In particular, diffusion processes conditioned to hit a specific point at a fixed future time, which are often referred to as diffusion bridges, are of great interest in situations where observations constrain the dynamics of a stochastic process. For example, in generative modeling, stochastic imputation between two given images, also known as the image translation task, uses diffusion bridges to model dynamics [Zhou et al., 2024, Zheng et al., 2024]. In the area of stochastic shape analysis and computational anatomy, random evolutions of biological shapes of organisms are modeled as non-linear diffusion bridges, and simulating such bridges is critical to solving inference and registration problems [Arnaudon et al., 2022, Baker et al., 2024a, Yang et al., 2024]. Finally, diffusion bridges play a crucial role in Bayesian inference and parameter estimations based on discrete-time observations. [Delyon and Hu, 2006, van der Meulen and Schauer, 2017, 2018, Pieschner and Fuchs, 2020]
@@
Simulation of diffusion bridges in either Euclidean spaces or manifolds is nontrivial since, in general, there is no closed-form expression for transition densities, which is key to constructing the conditioned dynamics via Doob’s $h$-transform [Rogers and Williams, 2000]. This task has gained a great deal of attention in the past decades [Beskos et al., 2006, Delyon and Hu, 2006, Schauer et al., 2017, Whitaker et al., 2016, Bierkens et al., 2021, Mider et al., 2021, Heng et al., 2022, Chau et al., 2024, Baker et al., 2024a]. Among them, one common approach is to use a proposed bridge process (called guided proposal) as an approximation to the true bridge. Then either MCMC or Sequential Monte Carlo (SMC) methods are deployed to sample the true bridge via the tractable likelihood ratio between the true and proposed bridges. Another solution is to use the score-matching technique [Hyvarinen, 2005, Vincent, 2011] to directly approximate the intractable score of the transition density using gradient-based optimization. Here, a neural network is trained with samples from the forward process [Heng et al., 2022] or adjoint process [Baker et al., 2024b], and plugged into numerical solving schemes, for example, Euler-Maruyama. Several recent studies deal with the extension of bridge simulation techniques beyond Euclidean spaces to manifolds [Sommer et al., 2017, Jensen and Sommer, 2023, Grong et al., 2024, Corstanje et al., 2024]. All of these rely on either a type of guided proposal or score matching.
@@
Both guided-proposal-based and score-learning-based bridge simulation methods have certain limitations: the guided proposal requires a careful choice of a certain “auxiliary process”. Mider et al. [2021] provided various strategies, but it is fair to say that guided proposals are mostly useful when combined with MCMC or Sequential Monte Carlo (SMC) methods. In case of a strongly nonlinear diffusion or high-dimensional diffusion, the simulation of bridges using guided proposals combined with MCMC (most notably the preconditioned Crank-Nicolson (pCN) scheme [Cotter et al., 2013]) or SMC may be computationally demanding. On the other hand, score-matching relies on sampling unconditioned processes, and it performs poorly for bridges conditioned on rare events, as the unconditioned process rarely explores those regions, resulting in inaccurate estimation. Additionally, the canonical score-matching loss requires the inversion of the matrix $\sigma\sigma^{T}$ where $\sigma$ is the diffusivity of the diffusion. This rules out hypo-elliptic diffusions and poses significant computational challenges for high-dimensional diffusions, where this matrix can become close to singularity, further exacerbating the difficulty of obtaining stable and accurate minimization of the loss.
@@
To address these issues, we introduce a new bridge simulation method called the neural guided diffusion bridge. It consists of the guided proposal introduced in Schauer et al. [2017] with an additional superimposed drift term that is parametrized by a neural network. The family of laws on path space induced by such proposals provides a rich variational family for approximating the law of the diffusion bridge. Once the variational approximation has been learned, independent samples can be generated at a cost similar to that of sampling the unconditional (forward) process. The contributions of this paper are as follows:
@@
  • We propose a simple diffusion bridge simulation method inspired by the guided proposal framework, avoiding the need for reverse-process modelling or intensive MCMC or SMC updates. Once the network has been trained, obtaining independent samples from the variational approximation is trivial and computationally cheap;
@@
  • Unlike score-learning-based simulation methods, which rely on unconditional samples for learning, our method is grounded to learn directly from conditional samples. This results in greater training efficiency.
@@
  • We validate the method through numerical experiments ranging from one-dimensional linear to high-dimensional nonlinear cases, offering qualitative and quantitative analyses. Advantages and disadvantages compared to the guided proposal [Mider et al., 2021] and two score-learning-based methods, Heng et al. [2022] and Baker et al. [2024b], are included.
@@ 2 Related work
Diffusion bridge simulation: This topic has received considerable attention over the past two decades and it is hard to give a short complete overview. Early contributions are Clark [1990], Chib et al. [2004], Delyon and Hu [2006], Beskos et al. [2006], Lin et al. [2010], and Golightly and Wilkinson [2010]. The approach of guided proposals that we use here was introduced in Schauer et al. [2017] for fully observed uniformly elliptic diffusions and later extend to partially observed hypo-elliptic diffusions in Bierkens et al. [2020].
@@
The method we propose is a combination of existing ideas. It used the guided proposals from Schauer et al. [2017] to construct a conditioned process, but learns an additional drift term parametrized by a neural net using variational inference.
Diffusion Schrödinger bridge: The diffusion bridge problem addressed in this paper may appear similar to the diffusion Schrödinger bridge (DSB) problem due to their names, but they are fundamentally different. A diffusion bridge refers to a diffusion process conditioned to start at one fixed point and reach another fixed point, while the DSB involves connecting two fixed marginal distributions and is often framed as an entropy-regularized optimal transport problem. Although DSB has gained attention for applications in generative modelling [Thornton et al., 2022, De Bortoli et al., 2021, Shi et al., 2024, Tang et al., 2024], it is important to recognize the distinctions between these problems.
Neural SDE: Neural SDEs extend neural ODEs [Chen et al., 2018] by incorporating inherent stochasticity, making them suitable for modelling data with stochastic dynamics. Research on neural SDEs can be broadly categorized into two areas: (1) modelling terminal state data [Tzen and Raginsky, 2019a, b], and (2) modelling entire data trajectories [Li et al., 2020, Kidger et al., 2021]. Our method aligns with the second category, as it employs trainable drift terms and incorporates end-point constraints to model the full data trajectory.
@@ 3.1 Problem statement
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space with filtration $\{\mathcal{F}_{t}\}_{t\in[0,T]}$ and $W$ be a $d_{w}$-dimensional $\mathbb{P}$-Wiener process. A $d$-dimensional $\{\mathcal{F}_{t}\}$-adapted diffusion process $X$ with the law of $\mathbb{P}$ is defined as the strong solution to the stochastic differential equation (SDE):
@@
The coefficient terms $b,\sigma$ are assumed to be Lipschitz continuous and of linear growth to guarantee the existence of a strong solution $X_{t}$ [Øksendal, 2014, Chapter 5.2]. In addition, we impose the standing assumption that $X$ admits smooth transition densities $p$ with respect to the Lebesgue measure $\lambda$ on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$, where $\mathcal{B}(\mathbb{R}^{d})$ is the Borel algebra of $\mathbb{R}^{d}$. That is, $\mathbb{P}(X_{t}\in A\mid X_{s}=x)=\int_{A}p(t,y\mid s,x)\lambda(\mathrm{d}y)$ for $0\leq s<t\leq T$, $A\subset\mathbb{R}^{d}$.
@@
Let $\mathbb{P}^{\circ},\mathbb{P}^{\star}$ and $\mathbb{P}^{\bullet}$ be measures on $(\Omega,\mathcal{F})$, we denote the laws of $X$ on $\mathcal{C}([0,T],\mathbb{R}^{d})$ under $\mathbb{P}^{\circ},\mathbb{P}^{\star}$ and $\mathbb{P}^{\bullet}$ by $\mathbb{L}^{\circ},\mathbb{L}^{\star}$ and $\mathbb{L}^{\bullet}$ respectively. For notational ease, the expectations under $\mathbb{P}^{\circ},\mathbb{P}^{\star}$ and $\mathbb{P}^{\bullet}$ (and similarly $\mathbb{L}^{\circ},\mathbb{L}^{\star}$ and $\mathbb{L}^{\bullet}$) are denoted by $\mathbb{E}^{\circ},\mathbb{E}^{\star}$ and $\mathbb{E}^{\bullet}$ respectively. The process $X$ under $\mathbb{L}^{\circ}$ and $\mathbb{L}^{\bullet}$ is sometimes denoted by $X^{\circ}$ and $X^{\bullet}$ respectively. For any measure $\mathbb{Q}$ on $(\Omega,\mathcal{F})$, we always denote its restriction to $\mathcal{F}_{t}$ by $\mathbb{Q}_{t}$.
@@
$\mathrm{d}X_{t}=\{b(t,X_{t})+a(t,X_{t})r(t,X_{t})\}\mathrm{d}t+\sigma(s,X_{s})\mathrm{d}W^{\star}_{s},\quad X_{0}=x_{0}.$ (2)
where $r(s,x)=\nabla_{x}\log h(s,x)$, $a(s,x)=\sigma(s,x)\sigma^{T}(s,x)$ and $W^{\star}$ is a $\mathbb{P}^{\star}$-Wiener process.
@@
A Bayesian interpretation of this result is obtained by assigning the endpoint $y$ a prior density $\pi(y)=p(T,y\mid 0,x_{0})$, and where the observation is given by $v\in\mathbb{R}^{d^{\prime}}$. The likelihood of this observation is $\ell(y\mid v)=q(v\mid y)$ and therefore $\xi(\mathrm{d}y)$ gives the posterior measure of $y$, conditioned on observing $v$. Therefore, under $\mathbb{P}^{\star}$, the process $X$ is constructed by first sampling the endpoint $y$ conditioned on the observation $v$ by $\ell(y\mid v)$, followed by sampling the bridge to $y$. In this paper, we will apply this result when $q(v\mid y)=\psi(v;Ly,\Sigma)$, where $L\in\mathbb{R}^{d\times d^{\prime}}$ with $d^{\prime}\leq d$ and $L$ is assumed to be of full (row) rank. Here, $\psi(x;\mu,\Sigma)$ denotes the density of the $\mathcal{N}(\mu,\Sigma)$-distribution, evaluated at $x$. For example, for a two-dimensional diffusion observing only the first component with small error corresponds to $L=\begin{bmatrix}1&0\end{bmatrix}$ and $\Sigma=\epsilon^{2}\mathbf{I}$, taking $\epsilon$ very small.
If $p$ were known in closed form, then the conditioned process could be directly sampled from Equation (2). This is rarely the case. For this reason, let $\tilde{X}$ be an auxiliary diffusion process that admits transition densities $\tilde{p}$ in closed form. Let $\tilde{h}(t,x)=\int_{\mathbb{R}^{d}}\tilde{p}(T,y\mid t,x)q(v\mid y)\nu(\mathrm{d}y)$. Define
@@
where $\mathcal{A}$ is the infinitesimal generator of the process $\tilde{X}$, i.e. for any $f$ in its domain $\mathcal{A}f(x)=\sum_{i}b_{i}(t,x)\partial_{i}f(t,x)+\frac{1}{2}\sum_{i,j}a_{i,j}(t,x)\partial_{ij}f(t,x)$. Under weak conditions (see e.g. Palmowski and Rolski [2002, Lemma 3.1]), $E_{t}$ is a mean-one martingale. If we define the change of measure $\mathrm{d}\mathbb{P}^{\circ}_{t}=E_{t}\mathrm{d}\mathbb{P}_{t}$, then, under $\mathbb{P}^{\circ}_{t}$, the process $X$ solves the SDE
$\mathrm{d}X_{t}=\{b(t,X_{t})+a(t,X_{t})\tilde{r}(t,X_{t})\}\mathrm{d}t+\sigma(t,X_{t})\mathrm{d}W^{\circ}_{t},\quad X_{0}=x_{0},$ (6)
where $\tilde{r}(t,x)=\nabla_{x}\log\tilde{h}(t,x)$ and $W^{\circ}$ is a $\mathbb{P}^{\circ}$-Wiener process. The process specified by the dynamics in Equation (6) is called the guided proposal, which is a process constructed to resemble the true conditioned process by replacing $r$ by $\tilde{r}$. Crucially, as its drift and diffusion coefficients are known in closed form, the guided proposal can be sampled using efficient numerical SDE solvers.
In Schauer et al. [2017] and Bierkens et al. [2020], precise conditions are given under which $\mathbb{P}^{\star}_{T}\ll\mathbb{P}^{\circ}_{T}$. In the case of conditioning on the event $\{LX_{T}=v\}$ –so there is no noise on the observation– this is subtle. We postpone a short discussion on this to Section 3.3. In all numerical examples considered in Section 5, we have ensured sufficient conditions are satisfied. This guarantees good behaviour when the noise level of the observation noise tends to zero. We then get the following theorem from Bierkens et al. [2020, Theorem 2.6] that states the change of laws from $\mathbb{L}^{\circ}$ to $\mathbb{L}^{\star}$.

### v3

@@ Abstract
We propose a novel method for simulating conditioned diffusion processes (diffusion bridges) in Euclidean spaces. By training a neural network to approximate bridge dynamics, our approach eliminates the need for computationally intensive Markov Chain Monte Carlo (MCMC) methods or score modeling. Compared to existing methods, it offers greater robustness across various diffusion specifications and conditioning scenarios. This applies in particular to rare events and multimodal distributions, which pose challenges for score-learning- and MCMC-based approaches. We introduce a flexible variational family, partially specified by a neural network, for approximating the diffusion bridge path measure. Once trained, it enables efficient sampling of independent bridges at a cost comparable to sampling the unconditioned (forward) process.
@@ 1 Introduction
Diffusion processes play a fundamental role in various fields such as mathematics, physics, evolutionary biology, and, recently, generative models. In particular, diffusion processes conditioned to hit a specific point at a fixed future time, which are often referred to as diffusion bridges, are of great interest in situations where observations constrain the dynamics of a stochastic process. For example, in generative modeling, stochastic imputation between two given images, also known as the image translation task, uses diffusion bridges to model dynamics [Zhou et al., 2024, Zheng et al., 2025]. In the area of stochastic shape analysis and computational anatomy, random evolutions of biological shapes of organisms are modeled as non-linear diffusion bridges, and simulating such bridges is critical to solving inference and registration problems [Arnaudon et al., 2022, Baker et al., 2024a, Yang et al., 2025]. Additionally, diffusion bridges play a crucial role in Bayesian inference and parameter estimations based on discrete-time observations. [Delyon and Hu, 2006, van der Meulen and Schauer, 2017, 2018, Pieschner and Fuchs, 2020]
@@
Simulation of diffusion bridges in either Euclidean space or manifolds is nontrivial since, in general, there is no closed-form expression for transition densities, which is key to constructing the conditioned dynamics via Doob’s $h$-transform [Rogers and Williams, 2000]. This task has gained a great deal of attention in the past decades [Beskos et al., 2006, Delyon and Hu, 2006, Schauer et al., 2017, Whitaker et al., 2016, Bierkens et al., 2021, Mider et al., 2021, Heng et al., 2022, Chau et al., 2024, Baker et al., 2024b]. Among them, one common approach is to use a proposed bridge process (called guided proposal) as an approximation to the true bridge. Then either MCMC or Sequential Monte Carlo (SMC) methods are deployed to sample the true bridge via the tractable likelihood ratio between the true and proposed bridges. Another solution is to use the score-matching technique [Hyvarinen, 2005, Vincent, 2011] to directly approximate the intractable score of the transition probability using gradient-based optimization. Here, a neural network is trained with samples from the unconditioned process [Heng et al., 2022] or adjoint process [Baker et al., 2024b], and plugged into numerical solving schemes, for example, Euler-Maruyama. Several recent studies deal with the extension of bridge simulation techniques beyond Euclidean spaces to manifolds [Sommer et al., 2017, Jensen and Sommer, 2023, Grong et al., 2024, Corstanje et al., 2024]. All of these rely on either a type of guided proposal or score matching.
@@
Both guided-proposal-based and score-learning-based bridge simulation methods have certain limitations: the guided proposal requires a careful choice of a certain “auxiliary process”. Mider et al. [2021] provided various strategies, but it is fair to say that guided proposals are mostly useful when combined with MCMC or Sequential Monte Carlo (SMC) methods. In case of a strongly nonlinear diffusion or high-dimensional diffusion, the simulation of bridges using guided proposals combined with MCMC (most notably the preconditioned Crank-Nicolson (pCN) scheme [Cotter et al., 2013]) or SMC may be computationally demanding. On the other hand, score-matching relies on sampling unconditioned processes, and it performs poorly for bridges conditioned on rare events, as the unconditioned process rarely explores those regions, resulting in inaccurate estimation. Additionally, the canonical score-matching loss requires the inversion of $\sigma\sigma^{\top}$ with $\sigma$ denoting the diffusion coefficient of the process. This rules out hypo-elliptic diffusions, where $\sigma\sigma^{\top}$ is singular. It also poses computational challenges for high-dimensional diffusions, further exacerbating the difficulty of obtaining stable and accurate minimization of the loss.
@@
To address these issues, we introduce a new bridge simulation method called neural guided diffusion bridge. It consists of the guided proposal introduced in Schauer et al. [2017] with an additional correction drift term that is parametrized by a learnable neural network. The family of laws on path space induced by such improved proposals provides a rich variational family for approximating the law of the diffusion bridge. Once the variational approximation has been learned, independent samples can be generated at a cost similar to that of sampling the unconditioned (forward) process. The contributions of this paper are as follows:
@@
  • We propose a simple diffusion bridge simulation method inspired by the guided proposal framework, avoiding the need for score modelling or intensive MCMC or SMC updates. Once the network has been trained, obtaining independent samples from the variational approximation is trivial and computationally cheap;
@@
  • Unlike score-learning-based simulation methods, which rely on unconditional samples for learning, our method is grounded to learn directly from conditional samples. This results in greater training efficiency, especially for learning the bridges that are conditioned on rare events.
@@
  • We validate the method through numerical experiments ranging from one-dimensional linear to high-dimensional nonlinear cases, offering qualitative and quantitative analyses. Advantages and disadvantages compared to the guided proposal Mider et al. [2021] and two score-learning-based methods, Heng et al. [2022] and Baker et al. [2024b], are included.
@@ 2 Related work
Diffusion bridge simulation: This topic has received considerable attention over the past two decades and it is hard to give a short complete overview. Early contributions are Clark [1990], Chib et al. [2004], Delyon and Hu [2006], Beskos et al. [2006], Lin et al. [2010], Golightly and Wilkinson [2010]. The approach of guided proposals that we use here was introduced in Schauer et al. [2017] for fully observed uniformly elliptic diffusions and later extend to partially observed hypo-elliptic diffusions in Bierkens et al. [2020].
@@
The method we propose is a combination of existing ideas. It used the guided proposals from Schauer et al. [2017] to construct a conditioned process, but learns an additional drift term parametrized by a neural network using variational inference.
Diffusion Schrödinger bridge: The diffusion bridge problem addressed in this paper may appear similar to the diffusion Schrödinger bridge (DSB) problem due to their names, but they are fundamentally different. A diffusion bridge is a process conditioned to start and end at specific points, often used for simulating individual sample paths under endpoint constraints. In contrast, a Schrödinger bridge connects two marginal distributions over time by finding the most likely stochastic process (relative to a reference diffusion) that matches these marginals. While both modify the original dynamics, the diffusion bridge imposes hard constraints on endpoints, whereas the Schrödinger bridge enforces them in distribution. Although DSB has gained attention for applications in generative modelling [Thornton et al., 2022, De Bortoli et al., 2021, Shi et al., 2024, Tang et al., 2024], it is important to recognize the distinctions between these problems.
Neural SDE: Neural SDEs generalize neural ODEs [Chen et al., 2018] by introducing stochasticity, enabling the modeling of systems with inherently random dynamics. Research in this area can be broadly divided into two categories: (1) modeling terminal state data [Tzen and Raginsky, 2019a, b], and (2) modeling entire trajectories [Li et al., 2020, Kidger et al., 2021]. Our approach falls into the latter category, leveraging trainable drift terms and end-point constraints to capture full trajectory dynamics.
@@ 3.1 Problem Statement
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space with filtration $\{\mathcal{F}_{t}\}_{t\in[0,T]}$, $W$ a $d_{w}$-dimensional $\mathbb{P}$-Wiener process, $b:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d\times d_{w}}$ the drift- and diffusion coefficients. A $d$-dimensional $\{\mathcal{F}_{t}\}$-adapted diffusion process $X$ with the law of $\mathbb{P}$ is defined as the strong solution to the stochastic differential equation (SDE):
@@
The coefficients $b,\sigma$ are assumed to be Lipschitz continuous and of linear growth to guarantee the existence of a strong solution $X_{t}$ [Øksendal, 2014, Chapter 5.2]. In addition, we impose the standing assumption that $X$ admits smooth transition densities $p$ with respect to the Lebesgue measure $\lambda$ on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$, where $\mathcal{B}(\mathbb{R}^{d})$ is the Borel algebra of $\mathbb{R}^{d}$. That is, $\mathbb{P}(X_{t}\in A\mid X_{s}=x)=\int_{A}p(t,y\mid s,x)\lambda(\mathrm{d}y)$ for $0\leq s<t\leq T$, $A\subset\mathbb{R}^{d}$.
@@
Let $\mathbb{P}^{\circ},\mathbb{P}^{\star}$ and $\mathbb{P}^{\bullet}$ be measures on $(\Omega,\mathcal{F})$, we denote the laws of $X$ on $\mathcal{C}([0,T],\mathbb{R}^{d})$ under $\mathbb{P}^{\circ},\mathbb{P}^{\star}$ and $\mathbb{P}^{\bullet}$ by $\mathbb{L}^{\circ},\mathbb{L}^{\star}$ and $\mathbb{L}^{\bullet}$ respectively. For notational ease, the expectations under $\mathbb{P}^{\circ},\mathbb{P}^{\star}$ and $\mathbb{P}^{\bullet}$ (and similarly $\mathbb{L}^{\circ},\mathbb{L}^{\star}$ and $\mathbb{L}^{\bullet}$) are denoted by $\mathbb{E}^{\circ},\mathbb{E}^{\star}$ and $\mathbb{E}^{\bullet}$ respectively. The process $X$ under $\mathbb{L}^{\circ},\mathbb{L}^{\star}$ and $\mathbb{L}^{\bullet}$ is sometimes denoted by $X^{\circ},X^{\star}$ and $X^{\bullet}$ respectively. For any measure $\mathbb{Q}$ on $(\Omega,\mathcal{F})$, we always denote its restriction to $\mathcal{F}_{t}$ by $\mathbb{Q}_{t}$.
@@
$\mathrm{d}X_{t}=\{b(t,X_{t})+a(t,X_{t})\,r(t,X_{t})\}\,\mathrm{d}t+\sigma(t,X_{t})\,\mathrm{d}W^{\star}_{s},\quad X_{0}=x_{0}.$ (2)
where $r(s,x)=\nabla_{x}\log h(s,x)$, $a(s,x)=\sigma(s,x)\sigma^{\top}(s,x)$ and $W^{\star}$ is a $\mathbb{P}^{\star}$-Wiener process.
@@
Remark 1.
A Bayesian interpretation of Equation 4 can be obtained by considering the following hierarchical model:
$v\mid y$ $\sim q(v\mid y),$ (5a)
$y$ $\sim p(T,y\mid 0,x_{0}).$ (5b)
Here, $y$ is considered as the parameter that gets assigned the prior density $p(T,y\mid 0,x_{0})$ and $v$ is the observation. Therefore, $\xi$ is the posterior distribution of $y$ and (3) shows that sampling of the conditioned process can be done by first sampling the endpoint $x_{T}$ from distribution $\xi$, followed by sampling a bridge connecting $x_{0}$ and $x_{T}$.
Throughout the paper, we will consider $q(v\mid y)=\psi(v;Ly,\Sigma)$, where $L\in\mathbb{R}^{d^{\prime}\times d}$ with $d^{\prime}\leq d$ and $L$ of full (row) rank. Here, $\psi(x;\mu,\Sigma)$ denotes the density of the $\mathcal{N}(\mu,\Sigma)$-distribution, evaluated at $x$. For example, for a two-dimensional diffusion $y=\begin{bmatrix}y_{1}&y_{2}\end{bmatrix}^{\top}$ observing only the first component $y_{1}$ corresponds to $L=\begin{bmatrix}1&0\end{bmatrix}$ as $Ly=y_{1}$. In our simulation experiments, we will assume $\Sigma=\epsilon^{2}\mathbf{I}$, for a small value of $\epsilon$, which is close to observing without error. Taking $\epsilon$ strictly positive stabilizes numerical computations.
3.2 Guided Proposal
If $p$ were known in closed form, then the conditioned process could be directly sampled from Equation 2. This is rarely the case. For this reason, let $\tilde{X}$ be an auxiliary diffusion process that admits transition densities $\tilde{p}$ in closed form. Let $\tilde{h}(t,x)=\int_{\mathbb{R}^{d}}\tilde{p}(T,y\mid t,x)q(v\mid y)\nu(\mathrm{d}y)$. Define
@@ 3.2 Guided Proposal
where $\mathcal{A}$ is the infinitesimal generator of the process $X$, i.e. for any $f$ in its domain $\mathcal{A}f(x)=\sum_{i}b_{i}(t,x)\partial_{i}f(t,x)+\frac{1}{2}\sum_{i,j}a_{ij}(t,x)\partial_{ij}f(t,x)$. Let $t<T$. Under weak conditions (see e.g. [Palmowski and Rolski, 2002, Lemma 3.1]), $\mathbb{E}[E_{t}]=1$. Using $\tilde{h}$ we can define the guided proposal.
Definition 1.
[Schauer et al., 2017] If we define the change of measure $\mathrm{d}\mathbb{P}^{\circ}_{t}=E_{t}\mathrm{d}\mathbb{P}_{t}$, then, under $\mathbb{P}^{\circ}_{t}$, the process $X$ solves the SDE
$\mathrm{d}X_{s}=\{b(s,X_{s})+a(s,X_{s})\tilde{r}(s,X_{s})\}\,\mathrm{d}s+\sigma(s,X_{s})\,\mathrm{d}W^{\circ}_{s},\quad X_{0}=x_{0}$ (7)
where $s\in[0,t]$, $\tilde{r}(t,x)=\nabla_{x}\log\tilde{h}(t,x)$ and $W^{\circ}$ is a $\mathbb{P}^{\circ}_{t}$-Wiener process. The process $X$ under the law $\mathbb{P}^{\circ}_{t}$ is known as the guided proposal.
Intuitively, $X^{\circ}$ is constructed to resemble the true conditioned process $X^{\star}$ by replacing $r$ by $\tilde{r}$. Crucially, as its drift and diffusion coefficients are known in closed form, the guided proposal can be sampled using efficient numerical SDE solvers such as Euler-Maruyama.
The definition of the guided process can be extended to $[0,T]$ by continuity. Whereas $\mathbb{P}_{t}\ll\mathbb{P}^{\circ}_{t}$ for $t<T$ the measures will typically be singular in the limit $t\uparrow T$. Nevertheless, $\mathbb{P}^{\star}_{t}\ll\mathbb{P}^{\circ}_{t}$ may still hold under this limiting operation and this is what matters for our purposes. In Schauer et al. [2017] and Bierkens et al. [2020], precise conditions are given under which $\mathbb{P}^{\star}_{T}\ll\mathbb{P}^{\circ}_{T}$. In the case of conditioning on the event $\{LX_{T}=v\}$ –so there is no noise on the observation– this is subtle. We postpone a short discussion on this to Section 3.4 to argue that all numerical examples considered in Section 5 will not break down in case the noise level on the observation, $\epsilon$, tends to zero. We then get the following theorem from [Bierkens et al., 2020, Theorem 2.6] that states the change of laws from $\mathbb{L}^{\circ}$ to $\mathbb{L}^{\star}$.